Bay Beach Industries wants to maintain their capital structure of 40% debt and 60% equity. The firm's tax rate is 34%. The firm can issue the following securities to finance the investments:

Bonds: Mortgage bonds can be issued at a pre-tax cost of 9 percent. Debentures can be issued at a pre-tax cost of 10.5 percent.

Common Equity: Some retained earnings will be available for investment. In addition, new common stock can be issued at the market price of $46. Flotation costs will be $3 per share. The recent common stock dividend was $3.60. Dividends are expected to grow at 6% in the future.

What is the cost of capital using mortgage bonds and internal equity?
Answer #1

kIE = [{D0 x (1 + g)} / P0] + g

= [($3.60 x 1.06) / $46] + 0.06 = 0.0830 + 0.06 = 0.1430, or 14.30%

After-tax kD = Pre-tax kD x (1 - t) = 9% x (1 - 0.34) = 5.94%

WACC = [wD x After-tax kD] + [wE x kE]

= [0.4 x 5.94%] + [0.6 x 14.30%] = 2.38% + 8.58% = 10.95%